OpenStudy (anonymous):

he boards that are stacked in the garage are each 0.0625 meters wide. So each board is ____________________ of a meter wide. A. \frac{1}{16} B. \frac{625}{100} C. \frac{1}{36} D. \frac{1}{625}

OpenStudy (anonymous):

@Nnesha

OpenStudy (fibonaccichick666):

can you name the places after the decimal?

OpenStudy (fibonaccichick666):

what place is the 5 in?

OpenStudy (tkhunny):

Have you considered performing the division and checking it out? B: 625/100 = 6.25 -- Nope, that's nowhere near 0.0625

OpenStudy (anonymous):

Sorry bacl

OpenStudy (anonymous):

I don't quite understand

OpenStudy (anonymous):

Anyone...?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=.0625

OpenStudy (anonymous):

\[.0625=\frac{625}{10000}\] which by some miracle reduces to \(\frac{1}{16}\)

OpenStudy (fibonaccichick666):

Do you know the place names? ex. ones place, tens, tenths, etc

OpenStudy (anonymous):

So A?

OpenStudy (tkhunny):

1/36 according to your calculator, is 0.027777 -- Nope. That's not 0.0625

OpenStudy (anonymous):

And yea I do

OpenStudy (anonymous):

^Wut

OpenStudy (fibonaccichick666):

well, if you know that, then you can create a fraction

OpenStudy (fibonaccichick666):

so what place is the 5 in?

OpenStudy (anonymous):

Thousandsths. WOuldn't that be 625/1000 than?

OpenStudy (fibonaccichick666):

I don't think it is in the thousandths

OpenStudy (fibonaccichick666):

count again

OpenStudy (anonymous):

O fail

OpenStudy (fibonaccichick666):

but if it was, then yes, it would be over 1000

OpenStudy (anonymous):

Ten Thousandths

OpenStudy (fibonaccichick666):

so which place is it, and what is the fraction then?

OpenStudy (anonymous):

625/10000

OpenStudy (fibonaccichick666):

there we go

OpenStudy (fibonaccichick666):

now, what number do you know will divide both the top and the bottom evenly?

OpenStudy (fibonaccichick666):

so that you can reduce it

OpenStudy (anonymous):

I don't know. I know how to do the problem. But I don't know the number

OpenStudy (fibonaccichick666):

well, what number will always divide a number that ends in 5 or 0?

OpenStudy (anonymous):

10? 1? 2.5?

OpenStudy (fibonaccichick666):

well, ten won't divide 5 right?

OpenStudy (fibonaccichick666):

1 yes, but that doesn't get us anywhere

OpenStudy (fibonaccichick666):

and 2.5 isn't a whole number

OpenStudy (fibonaccichick666):

if I asked you what the multiples of 5 are, can you please tell me them?

OpenStudy (anonymous):

I know them. but what exactly are you asking for? it's a long list of 5s

OpenStudy (fibonaccichick666):

just give me some, in order

OpenStudy (anonymous):

5 10 15 20 25 30 35 40 45 50 55 60. (Hoping that was what you were askinjg for)

OpenStudy (fibonaccichick666):

that's fine, now what do you notice about the last digit?

OpenStudy (anonymous):

it's a 0

OpenStudy (anonymous):

or a 5 lol

OpenStudy (fibonaccichick666):

there we go, so now, what number do you now will divide both the top and bottom of \[\frac{625}{10000}\] in order to reduce the fraction?

OpenStudy (anonymous):

25 or 50?

OpenStudy (fibonaccichick666):

make it easier

OpenStudy (fibonaccichick666):

pick a number you know will always divide the two

OpenStudy (anonymous):

25

OpenStudy (fibonaccichick666):

hint, why would I have made you list those multiples

OpenStudy (anonymous):

I am so dumb right now. Sorry

OpenStudy (fibonaccichick666):

it's ok, we all have our moments, although 25 does work, I really want you to see a way to reduce this when a big number doesn't pop out at you

OpenStudy (anonymous):

625 and 10000 can both be divided by 25

OpenStudy (anonymous):

unless 5 is the answer

OpenStudy (fibonaccichick666):

well, 5 is what I was looking for

OpenStudy (anonymous):

XD

OpenStudy (fibonaccichick666):

so if you do that division, what do you get?

OpenStudy (fibonaccichick666):

then we continue the simplifying game

OpenStudy (anonymous):

125/2000 I HATE SIMPLYIFIYING and I am very bad at it

OpenStudy (fibonaccichick666):

ok, now what number will divide both 125 and 2000?

OpenStudy (anonymous):

do i continue to divide by 5?

OpenStudy (fibonaccichick666):

yep, and what do you get

OpenStudy (anonymous):

1/16? unless I did something wrong

OpenStudy (anonymous):

^Prob did something wrong

OpenStudy (fibonaccichick666):

well, interestingly enough, it simplifies to 25/400 then 5/80 then . 1/16

OpenStudy (fibonaccichick666):

all by using 5's to make it smaller

OpenStudy (fibonaccichick666):

so, just remember, when simplifying, use small numbers. You may have a few more steps, but it is a lot easier. Most numbers will be divisible by 2,3,5,7,11 and that will be mostly all you'll need

OpenStudy (anonymous):

Can you help me with more other questions? I think I have about an hour left X:

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

make a new post, and I'll see what I can do

OpenStudy (anonymous):

ok thankjs!